1. If the periodic rate is 2%, and the interest rate is compounded quarterly, what is the APR?

2. If the APR is 11% with semi-annual compounding, what it the EFF or Effective Annual Rate (EAR)?

3. A lender offers to lend you $250,000. The loan calls for annual payments $25,000 for 20 years. What interest rate is the company charging you?
Answer #1
1) APR= Periodic rate * periods per year
APR = 2% * 4 quarter
=8%
2)
Effective Annual Rate =(1+r/n)^n -1
r= rate
n= number of periods per annuam
=(1+0.11/2)^2 -1
=(1+0.055)^2
=11.30%
3) Present Value Of An Annuity
= C*[1-(1+i)^-n]/i]
Where,
C= Cash Flow per period
i = interest rate per period
n=number of period
$250000= $25000[ 1-(1+r)^-20 /r]
10 = [ 1-(1+r)^-20 /r]
r = 0.07755
Interest rate = 7.76%